OpenStudy (anonymous):

Solve by factoring. 3a^2+5=-8a

OpenStudy (anonymous):

3a^2+8a+5 = 0 use abc formula

OpenStudy (anonymous):

3a^2 + 8a + 5 = 0 =>3a^2 + 5a + 3a + 5 = 0 => 3a(a+1) + 5 (a+1) = 0 => (3a+5) + (a+1) = 0 a = -1 or -5/3

OpenStudy (anonymous):

3a^2+8a+5=0 (3x+8)(x+1)

OpenStudy (anonymous):

sorry, the answer is: (3x+5)(x+1)